Kindly, tell me how this equation will be solved for distance "d", please.

(d/2)/12 + (d/2)/15 = 3

Thank you.
1
Expert's answer
2020-07-16T10:57:49-0400

Given,

"\dfrac{\frac{d}{2}}{12}+\dfrac{\frac{d}{2}}{15}=3"

Let, "p=\frac{d}{2}" ,thus new equation will be

"\frac{p}{12}+\frac{p}{15}=3"

Now,

"\text{LCM}(12,15)=60"

Thus,

"\frac{5p}{60}+\frac{4p}{60}=3\\\implies \frac{9p}{60}=3"

Multiply both side of the above equation by "60" we get

"9p=180"

Now divide both side of the above equation by "9" ,thus "p=20" but "p=\frac{d}{2}" ,hence

"d/2=20\implies d=40"